# Dynamic object goal pushing with mobile manipulators through model-free constrained reinforcement learning

### v1

@@ I Introduction
Moving and reorienting heavy or bulky objects along large and complex real-world pathways requires combining mobility and manipulation. This task is achievable through non-prehensile pushing actions without requiring a dedicated gripper or the need to grasp a handle on the object. In real-world scenarios, however, the object and terrain’s physical properties (e.g. mass, size, friction coefficient) are typically unknown and can reduce a controller’s performance. Additionally, non-prehensile pushing interaction may yield relative motion between the robot and object at the contact point, e.g. contact sliding or relative rotation. This motion necessitates a controller capable of reactively adapting the pushing location and direction by dynamically breaking and making contact at new locations with the object. We refer to this behavior as contact switching.
Achieving online contact switching behavior during non-prehensile pushing is challenging with model-based techniques [1, 2] or controllers that rely solely on force/tactile feedback [3, 4]. As a result, recent works leverage Reinforcement Learning (RL) to address contact switching [5] and demonstrate notable robustness against unknown objects [6]. Despite these achievements, the method in [5] is limited to fixed-base manipulators pushing a lightweight, small object on a table. Jeon et al. [6] achieve object pushing with the base of a mobile robot, without an arm. In both cases, the policies generate only 2D planar motion commands, which are insufficient for manipulating objects that are prone to toppling (e.g. objects with a thin base, large CoM height, or high friction coefficient flooring). To address this limitation, we focus on interacting with objects by pushing them to different 3D locations on their surface, enabling more versatile and stable manipulation.
@@
Our controller demonstrates robustness to unknown objects with different physical properties, such as mass, material, size, and shape. It autonomously handles contact switching by dynamically identifying the contact points and adjusting the push direction. Even though the policy only observes the object’s pose, it adapts to its xy-footprint and lowers the push point for thinner objects to maintain stability.
@@ IV-B Constraint satisfaction
On the hardware, actuation limits comprise the most challenging constraints to satisfy, in particular, the position limit for the shoulder flexion-extension joint (second arm joint). The pushing task requires the robot’s arm to reach low, especially for thin objects, while the position limit for this joint is slightly larger than 90 deg (with the zero on the positive axis of the base frame). In practice, this requires the controller to operate this joint very close to the limit during most of the task. Nevertheless, this was achieved during the extensive tests that we carried out.

### v2

@@ I Introduction
Moving and reorienting heavy or bulky objects along large and complex real-world pathways requires combining mobility and manipulation. This task is achievable through non-prehensile pushing actions without requiring a dedicated gripper or the need to grasp a handle on the object. In real-world scenarios, however, the object and terrain’s physical properties (e.g. mass, size, friction coefficient) are typically unknown and can reduce a controller’s performance. Additionally, non-prehensile pushing interaction may yield relative motion between the robot and object at the contact point, e.g. contact sliding or relative rotation. This motion necessitates a controller capable of reactively adapting the pushing location and direction by dynamically breaking and making contact at new locations with the object. We refer to this contact-rich behavior as contact switching.
Achieving online contact switching behavior during pushing is challenging with model-based techniques [1, 2] or controllers that rely solely on force/tactile feedback [3, 4]. As a result, recent works leverage Reinforcement Learning (RL) to address contact switching [5] and demonstrate notable robustness against unknown objects [6]. Despite these achievements, the method in [5] is limited to fixed-base manipulators pushing a lightweight, small object on a table. Jeon et al. [6] achieve object pushing with the base of a mobile robot, without an arm. In both cases, the policies generate only 2D planar motion commands, which are insufficient for manipulating objects that are prone to toppling (e.g. objects with a thin base, large CoM height, or high friction coefficient flooring). To address this limitation, we focus on interacting with objects by pushing them to different 3D locations on their surface, enabling more versatile and stable manipulation.
@@
Our controller demonstrates robustness to unknown objects with differing physical properties, such as mass, material, size, and shape. It autonomously handles contact switching by dynamically identifying the contact points and adjusting the push direction. Even though the policy only observes the object’s pose, it adapts to its xy-footprint and lowers the pushing location for thinner objects to maintain stability.
@@ IV-B Constraint satisfaction
On the hardware, actuation limits comprise the most challenging constraints to satisfy, in particular, the position limit for the shoulder flexion-extension joint (arm’s second joint). The pushing task requires the robot’s arm to reach low, especially for thin objects, while the position limit for this joint is slightly larger than 90 deg (with the zero on the positive axis of the base frame). In practice, this requires the controller to operate this joint very close to the limit during most of the task. Nevertheless, this was achieved during the extensive tests that we carried out.